I tried this one for at least 45 minutes, and really need help, please.
Mr. Elliot is extremely careful with his money. He recently invested part of his $150,000 life savings in gold bullion and the rest in a nuclear blast-proof bank. Mr. Elliot earned a 5% profit on his gold bullion and a 3% profit on the money in the bank. If he made a total of $5,500 from both investments, how much did he invest in gold bullion?
I tried both
.03x + .05(5,500 - x) = 150,000
and
.03x + .05(150,000 - x) + 5,500
solving for both, but neither came out right.
Thanks in advance for your help!

Answer by stanbon(75887)

Mr. Elliot is extremely careful with his money. He recently invested part of his $150,000 life savings in gold bullion and the rest in a nuclear blast-proof bank. Mr. Elliot earned a 5% profit on his gold bullion and a 3% profit on the money in the bank. If he made a total of $5,500 from both investments, how much did he invest in gold bullion?
I tried both
Equation:
Interest + interest = interest
.03x + .05(150,000 - x) = 5,500
----
Multiply thru by 100 to get:
3x + 5*150000 - 5x = 550000
----
-2x = 550000-750000
-2x = -200000
x = $100,000 (amt. invested at 3%
150,000-x = $50,000 (amt. invested at 5%
==================
Cheers,
Stan H.
==============
=====================